Evaluate T AB C− given that 2 2 0 1 2 3 3 1 , and 0 1 3 0 1 2 12 7 A B C       − − − = = =             


1
Expert's answer
2021-06-07T12:14:26-0400

Part 1

"\begin{pmatrix}2&2&0\\ 1&2&3\\ 3&1&0\end{pmatrix}^T" = "\begin{pmatrix}2&1&3\\ 2&2&1\\ 0&3&0\end{pmatrix}"

Part 2

"\begin{pmatrix}0&1&3\\ 0&1&2\\ 1&2&7\end{pmatrix}^T" "=\begin{pmatrix}0&0&1\\ 1&1&2\\ 3&2&7\end{pmatrix}"